# revtex4-2Repair the float

### v1

@@ V Results and discussion
As to the dephasing during the $\pi$-pulses, typical length $\Delta t_{\pi}$ of single $\pi$-pulses varied between 20 and 220 ns, with the lower limit $\Delta t_{\mathrm{min}}$ determined either by the qubit anharmonicity ($\Delta t_{\mathrm{min}}>h/|\alpha|\simeq 1/200\,$MHz $=5\,$ns) or by the dynamical range of instruments and total line attenuation of the setup (see Fig. 8). We present experimental data of the temperature dependence of the dephasing time $\tau_{\varphi}$ in Fig. 2 b. The times $\tau_{\varphi}^{R}$ and $\tau_{\varphi}^{E}$ are extracted from the measurements of Ramsey oscillations and Hahn echo measurements, respectively, as $\tau_{\varphi}^{R,E}=1/\gamma_{\varphi}^{R,E}=1/\left(\gamma_{2}^{R,E}-\gamma_{1}/2\right)$, where the decoherence rates $\gamma_{2}^{R,E}$ describe Ramsey oscillations and Hahn echo experiments. While in the low-temperature limit $\tau_{\varphi}^{R,E}\gtrapprox\tau_{1}$, with $\tau_{\varphi}^{R}$ exhibiting a clearly visible saturation below $50-70$ mK, the dephasing time starts decreasing at lower temperatures than $\tau_{1}$ and reaches $1-2\,\mu$s at 150 mK, becoming difficult to measure at $220-250\,$mK due to the finite ring-up time of the readout resonator $\sim Q_{ext}/\omega_{r}\simeq 1\,\mu$s. Similarly to the analysis of the relaxation time, we can estimate influence of quasiparticles on the dephasing, using the approach presented in Ref. [33]. The pure dephasing rate due to tunneling of equilibrium quasiparticles is given by expression
@@
Population ratios $A,\ B$ and $C$ can be extracted from the measurement data $x_{i},\ y_{j}$ by three different methods each, which we denote as $A_{1},A_{2},A_{3},B_{1},\dots,C_{3}$ (see Tab. 2). Qualitatively speaking, the difference between the methods is the order of the $\pi$-pulses in the control sequence. Because of the qubit relaxation and relatively low population of the $f$-level, the resulting signals can have different signal-to-noise (SNR) ratio, giving us some guidance on how to choose the best one (see Fig. 5 and discussion in Sec. VI and Appendix C).
An example of population measurements in terms of datasets $A_{1},A_{2},\dots C_{3}$ (see exact expressions in Appendix in Table 2) for device R4-I can be seen in Fig. 2 c. While according to expressions (7), (8), in the low-temperature limit $A\rightarrow 1$ and $B,C\rightarrow 0$, there is a clear saturation at $p_{e}/p_{g}\approx 0.024$, which agrees with $T_{\mathrm{eff}}\approx 85\,$mK, calculated by Eq. (2). Above 170 mK the population readings have increased deviation from the analytical expressions caused by the enhanced relaxation.
Measurements of the qubit effective temperature for devices R2-I, R4-I and Q2-III extracted from $A_{2}$ are shown in Fig. 2 d and demonstrate a linear behaviour in the temperature range $120-230$ mK close to one-to-one correspondence with the temperature of the mixing chamber stage $T_{\mathrm{MXC}}$ of the cryostat. At lower temperatures of the cryostat $T_{\mathrm{eff}}$ exhibits saturation, reaching $T_{\mathrm{eff}}^{0}\approx 65-85\,$mK, which coincides with the saturation temperature estimated from the population in Fig. 2 c. This elevated temperature compared to $T_{\mathrm{MXC}}$ is most probably caused by imperfect shielding against thermal photons and insufficient filtering of the measurement lines. For comparison, $T_{\mathrm{eff}}$ extracted using all nine methods $A_{1},A_{2},A_{3},B_{1},\dots,C_{3}$ are shown in Fig. 3 a. A thorough explanation of the deviations between the experimental data and the expected behaviour above 170 mK in Figs. 2 c, d requires consideration of the time evolution of the qubit population due to its finite lifetime, which we discuss in the following subsection.
V.1 Comparison of the Experimental Data with the Numerical Model of the Qubit Population Evolution
We begin with a qubit, which is thermalized at a temperature $T$, setting initial populations $p_{g,e,f}^{(0)}(T)$ and then imitate application of $\pi_{ge}$- and $\pi_{ef}$-pulses by instantaneous swapping of the corresponding level populations with finite efficiencies, designated as $\delta_{ge,ef}\in[0,1]$ (see Fig. 1 e). For example, a $\pi_{ge}$-pulse with $\delta_{ge}=0.9$ means that a qubit state with $p_{g}=1.0,\ p_{e}=0.0,\ p_{f}=0.0$ would turn result into $p_{g}=0.1,\ p_{e}=0.9,\ p_{f}=0.0$. Then we calculate the time evolution of the population during the readout time, and multiply it by arbitrary pure state responses $\varphi_{i}$. Next we construct the population ratios $A,\ B$ and $C$ and calculate the effective temperature (see Fig. 3 b). A detailed description of the model is provided in Appendix B.
The numerical model can qualitatively reproduce the deviation of the qubit population functions $A,B$ and $C$ from the analytical expressions (see Fig. 6 in App. B), as well as the resulting deviations of the temperatures $T_{\mathrm{eff}}$ extracted with different methods in the higher temperature range, which is presented in Fig. 3. This divergence is a consequence of the temperature-dependent qubit decay times $\tau_{1}^{ge}$ and $\tau_{1}^{ef}$, caused by quasiparticles. For the simulation results presented in Fig. 3 b, we use frequencies $\omega_{ge}$ and $\omega_{ef}$ and the temperature dependent relaxation time $\tau_{1}^{ge}(T)=2\pi/\left(\gamma_{1}^{qp}(T)+\gamma_{1}^{0}\right)$, which correspond to the sample R4-I shown as the red line in Fig. 2 b, and $\tau_{1}^{ef}\approx\tau_{1}^{ge}/2$, valid for a transmon qubit [40]. For ideal $\pi$-pulses, the effective temperatures $T_{A1},T_{A2},\dots,T_{C3}$ extracted from different ratios $A_{1},A_{2},\dots,C_{3}$, respectively, demonstrate almost one-to-one behaviour for the temperature range $30-200$ mK, when the qubit lifetimes are relatively long. But above 200 mK, when $\tau_{1}$ is significantly suppressed, there is an increasing deviation between these dependencies. Importantly, this does not indicate lack of thermalization of the qubit, rather than showing the limits of applicability of the given thermometry protocol. Moreover, in this temperature limit the values $T_{A1},T_{A2},\dots,T_{C3}$ start to depend on the pure state responses $\varphi_{i}$, which eliminates the main advantage of the protocol. In fact, Eqs. (III) are no longer valid, as the qubit population changes significantly during the readout pulse.
Another error in determining the effective temperature of the qubit is the finite efficiency of the pulses. As seen in Fig. 3 b, the divergence of $T_{A1},T_{A2},\dots,T_{C3}$ grows monotonously with the temperature and is noticeable already at $100-150$ mK, which is a characteristic of the protocol. The plot, showing the relative error between the extracted temperature and the qubit thermalization temperature averaged over different methods $A_{1},A_{2},\dots,C_{3}$ is shown in Fig. 7 in Appendix B.
Figure 3: Effective qubit temperature extracted from the population distribution measurements. a) Experimental data for the qubit R4-I for all nine ratios $A_{1},A_{2},A_{3},B_{1},\dots,C_{3}$. b) Effective qubit temperature, extracted from the qubit population distribution according to the numerical model (see details in main text). Using the device parameters of R4-I listed in Table 1, we initialize the qubit population $p_{i}(T),i\in\{g,e,f\}$ Boltzmann-distributed at temperature $T_{\mathrm{qubit}}$ and imitate the measurement procedure. We consider the temperature dependence of qubit lifetime $\tau_{1}$, which we approximate from experimental data shown in Fig. 2. Solid lines show the result of simulating perfect $\pi$ pulses, which we compare to finite pulse efficiency $\delta_{ge}=0.9,\delta_{ef}=0.8$ (dotted lines). The simulation was done with parameters $2\tau_{1}^{ef}(T_{\mathrm{qubit}}\rightarrow 0)=\tau_{1}^{ge}(T_{\mathrm{qubit}}\rightarrow 0)=5.5\,\mu s$, $\pi$-pulse duration is 165 ns and readout pulse length $\Delta t_{\mathrm{RO}}=2\,\mu$s.
V.2 Qubit Thermalization
In order to make sure, that the qubits used in our measurements thermalize to a single thermal bath, which is the substrate, we analyze the relaxation rate $\gamma_{1}$ and photon number $n\left(\omega_{ge},T\right)$ as functions of the qubit effective temperature $T_{\mathrm{eff}}$. For example, if a qubit is in a steady state with several ohmic baths, then from (3) and (25) we can write down the following temperature dependencies for $\gamma_{1}$ and $n\left(\omega_{ge},T_{\mathrm{eff}}\right)$:
$\gamma_{1}(T$ $)=\gamma_{1}^{0}\left[2\,n(\omega_{ge},T_{\mathrm{eff}})+1\right]=\gamma_{1}^{0}\,\coth{\frac{\hbar\omega_{ge}}{2k_{B}T_{\mathrm{eff}}}},$ (12)
@@ V.2 Qubit Thermalization
$n(\omega_{ge},T_{\mathrm{eff}})=\frac{\gamma_{1}^{0}}{\sum_{i}\gamma_{1}^{(i)}}n\left(\omega_{ge},T_{\mathrm{MXC}}\right)+n_{0},$ (13)
where we denote the relaxation rate and the residual qubit population at the base temperature as $\gamma_{1}^{0}$ and $n_{0}$, respectively. The corresponding experimental data is presented in Fig. 4. The photon number dependence of $\gamma_{1}$ (see Fig. 4 a) exhibits two regimes: the linear growth described by Eq. (12), where the derivative $\partial\gamma_{1}/\partial n$ is twice bigger than the offset $n_{0}$, and the sharp rise due to the quasiparticle relaxation.
The dependence between the photon number $n\left(\omega_{ge},T_{\mathrm{eff}}\right)$, which corresponds to the qubit effective temperature, and $n\left(\omega_{ge},T_{\mathrm{MXC}}\right)$, defined by the cryostat temperature, characterizes the general process of the qubit thermalization (see Fig. 4 b). In case of a single thermal bath coupled to the qubit, the dependence would be exactly one-to-one, whereas the presented data shows finite photon population $n_{0}$ in the low-temperature limit and the slope slightly less than unity, witnessing coupling to other thermal baths. One of those could be a residual radiation field, as well as a bath of TLSes or nonequilibrium quasiparticles with the effective temperature $65-85$ mK, which sets the lower limit for the population of the qubit at the lower temperatures of the cryostat.
Figure 4: a) Measurements of the qubit relaxation rate $\gamma_{1}$ as a function of the photon number $n\left(\omega_{ge},T_{\mathrm{eff}}\right)$ for qubits R2-I and Q2-III. The solid lines are linear fits $y=kx+b$, with parameters $k=0.71\pm 0.23$ MHz/photon and $b=0.41\pm 0.07$ MHz for R2-I and $k=0.63\pm 0.39$ MHz/photon and $b=0.37\pm 0.09$ MHz for Q2-III, respectively, in line with Eq. (12), i.e. $k\simeq 2b$. The dashed lines show the quasiparticle-induced relaxation $\gamma_{1}^{qp}$ according to Eq. (9). b) Dependence between the photon numbers $n\left(\omega_{ge},T_{\mathrm{eff}}\right)$ and $n\left(\omega_{ge},T_{\mathrm{MXC}}\right)$. The dashed line shows one-to-one correspondence and the solid lines are linear fits with parameters $k=0.965\pm 0.114$ and $b=0.017\pm 0.037$ for R2-I and $k=0.900\pm 0.140$ and $b=0.012\pm 0.045$ for Q2-III, respectively. Here, $k=\gamma_{1}^{0}/\sum_{i}\gamma_{1}^{(i)}\approx 1$ (see Eq. 13) means that thermalization of the qubit is dominated by its coupling to the thermal bath with the temperature $T_{\mathrm{MXC}}$.
VI Thermometer Limitations and Ways of Improvement
VI.1 Temperature Range
The Al-based transmon thermometers studied in this work showed a relatively narrow working temperature range $60-220$ mK. In this subsection we discuss possibilities of widening this regime.
A higher residual temperature of the qubits $T_{\mathrm{eff}}=65-85$ mK compared to the base cryostat temperature $T_{\mathrm{base}}\approx 20$ mK is most likely caused by both improper shielding of the sample stage and filtering of the measurement lines [43], likely due to nonequlibrium quasiparticles [27] or TLSes [28, 34]. Recent works demonstrated the possibility of achieving effective qubit temperatures of $30-45$ mK [27, 28, 30, 44].
When considering the higher temperature limit of the presented method the following aspects should be taken into account: i) consideration of only the three lowest states is valid until the population of the next $d$-level is negligible, otherwise, at higher temperatures, the system (III) should be supplemented with appropriate terms having $\varphi_{d}$ and $p_{d}$; ii) suppression of the qubit relaxation time $\tau_{1}$ due to growing quasiparticle density above 200 mK; iii) sufficiently long coherence time for the qubit state preparation. In agreement with the model, quasiparticle suppression of $\tau_{1}$ leads to a significant influence on the population measurements, which in turn affects adversely on determining the effective temperature. This is a hard limit due to the exponential growth of the quasiparticle density with temperature.
Another possibility is employment of materials with a larger superconducting gap to shift the quasiparticle suppression limit to higher temperatures. This might be done by using Nb-based tunnel junctions [45, 46, 47, 48] with the critical temperature $T_{c}^{\mathrm{Nb}}=7.6-8.7$ K or by replacing the tunnel junction with a nanowire made of granular Al having critical temperatures $T_{c}^{\mathrm{grAl}}$ up to 3.15 K [49, 50, 51], which is 2.5 times larger than critical temperature of bulk aluminium, $T_{c}^{\mathrm{Al,bulk}}=1.2$ K.
The qubit state preparation is also of big importance in the population distribution measurement presented here. In this sense, the qubit coherence time, or more strictly speaking, the dephasing time should be longer than the control pulse sequences $\left(\tau_{\varphi}\gtrsim 3\Delta t_{\pi}\right)$. Since the dephasing rate is growing with the temperature increase, it limits duration of the pulse sequences at higher temperatures.
VI.2 Signal-to-Noise Ratio and Accuracy
@@
The parameters of the experimental setup can be fine-tuned to improve both the temperature measurement accuracy and the range of the cryostat temperatures, so that it reaches the quasiparticle limit. First of all, improvement of the $\pi$-pulse efficiency affects positively on the measurement accuracy. Secondly, as mentioned above, the short lifetime of the qubit is the main restriction of the method. We can divide the qubit state decay into two parts: i) decay during the control sequence and ii) decay during the readout. The first process leads to change of the qubit population distribution $p_{i}$ and can be associated to imperfect control pulses, leading to increased deviation between $T_{A1},T_{A2},\dots,T_{C3}$. Decay during the readout pulse is defined mainly by the ring-up time of the resonator, and diminishes distinguishability between the qubit states (i.e. difference between $\varphi_{i}$ on the $I-Q$ plane), thus reducing the SNR, and also distorting the extracted effective temperature. The main method of improving the SNR is optimizing the ratio $\chi/\kappa$ for the readout resonator, where $\chi$ is the dispersive shift and $\kappa$ is the linewidth, as well as an appropriate readout frequency detuning for a given $\chi/\kappa$, which directly affects the pure state responses $\varphi_{i}$ [52].
@@
The minimal temperature measurement error is defined by the Cramer-Rao bound from the quantum Fisher information (QFI) [5], where we can use an expression for the relative error $\left|\Delta T/\langle T\rangle\right|_{\mathrm{sm}}$ of a single measurement for a two-level thermometer with energy separation $\hbar\omega_{ge}$ and $N$-fold degenerate excited level as
$\left(\frac{\Delta T}{\langle T\rangle}\right)^{2}_{\mathrm{sm}}=\frac{\left(N-1+e^{x_{ge}}\right)^{2}}{(N-1)x_{ge}^{2}e^{x_{ge}}},$ (14)
$x_{ge}=\hbar\omega_{ge}/k_{B}T$. An analogous expression can be derived for the three-level case considered in our work,
$\left(\frac{\Delta T}{\langle T\rangle}\right)^{2}_{\mathrm{sm}}=\frac{\left(e^{x_{ge}+x_{gf}}+e^{x_{ge}}+e^{x_{gf}}\right)^{2}}{\left[x_{ge}^{2}e^{x_{gf}}+x_{gf}^{2}e^{x_{ge}}+\left(x_{ge}+x_{gf}\right)^{2}\right]e^{x_{ge}+x_{gf}}},$ (15)
here $x_{gf}=\hbar\omega_{gf}/k_{B}T$ and $\omega_{gf}=2\omega_{ge}-|\alpha|$. Taking into account averaging over $N=2^{17}$ repetitions, one gets the lower bound for the relative error $\left(\Delta T/\langle T\rangle\right)^{2}=\left(\Delta T/\langle T\rangle\right)^{2}_{\mathrm{sm}}/N\approx 5.0\cdot 10^{-5}$ or $\left|\Delta T/\langle T\rangle\right|\approx 0.7\%$ for $\langle T\rangle=65$ mK and qubit frequency $\omega_{ge}=7.04$ GHz. The noise equivalent temperature $\mathrm{NET}=\sqrt{\left|\Delta T\right|^{2}\Delta t_{\mathrm{meas}}}$ with the measurement time $\Delta t_{\mathrm{meas}}=29$ s for our setup in the QFI-limited case would be $\mathrm{NET}_{\mathrm{QFI}}\approx 2.5$ mK$/\sqrt{\mathrm{Hz}}$. These estimates are an order of magnitude lower than the experimental errors. The data shown in Fig. 5 a has the relative errors $\left|\Delta T_{A}/\langle T_{A}\rangle\right|=14-42\%$, $\left|\Delta T_{B}/\langle T_{B}\rangle\right|=8.6-9.4\%$ and $\left|\Delta T_{C}/\langle T_{C}\rangle\right|=8.7-9.4\%$ at the base cryostat temperature $T_{\mathrm{MXC}}=22$ mK, which corresponds to $\mathrm{NET}\approx 28$ mK$/\sqrt{\mathrm{Hz}}$.
Figure 5: a) Time traces of the effective qubit temperature for the sample Q2-III at the base cryostat temperature $T_{\mathrm{MXC}}\approx 22\,$mK (shown as the black dashed line). b) Relative error of a single temperature measurement as a function of temperature. The symbols denote experimental time averaged data obtained from time traces of three samples: R2-I (green), R3-II (blue) and Q2-III (red). The error bars show the standard deviation. The curves denote the minimal error (i.e. the red filled area cannot be realized), defined by QFI for a two-level system (red dashed line) from the Eq. (14) for $N=2$, for a three level system with anharmonicity (solid dark red line) using Eq. (15), and for a $N$-level system with $N-1$ times degenerate excited level (red dotted line), calculated using Eq. (14) with $N=10$. The energy separations correspond to those defined for the sample Q2-III. The black dotted line shows the relative error for a three-level transmon with $\omega_{ge}/2\pi=3.00\,$GHz and $\omega_{ef}/2\pi=2.763$ GHz with the exponential rise shifted towards lower temperatures by $10-15$ mK.
Notably, if the thermometer has a constant level separation, the relative error grows exponentially at $T\ll\hbar\omega_{ge}/k_{B}$ (see Fig. 5), which is, however, not observed experimentally in our case, as the qubit effective temperature saturates at $T_{\mathrm{eff}}^{0}\approx 65-85$ mK. Our system is a three-level quantum thermometer with a constant anharmonic spectrum, which refers to sub-optimal quantum thermometers [5]. Knowing the dependence of QFI on the spectrum of the system, it is possible to optimize it for a better performance in the experimentally realistic temperature range by choosing an appropriate level separation and anharmonicity.
VI.3 Operation Speed

### v3

@@ V.1.2 Suppression of Coherence
The typical length $\Delta t_{\pi}$ of separate $\pi$-pulses varied between 60 and 220 ns, with the lower limit $\Delta t_{\mathrm{min}}$ determined either by the qubit anharmonicity ($\Delta t_{\mathrm{min}}>h/|\alpha|\simeq 1/200\,$MHz $=5\,$ns) or by the dynamical range of the instruments and the total attenuation of the setup (see Fig. 9). We present experimental data of the temperature dependence of the dephasing time $\tau_{\varphi}$ in Fig. 5 b. The times $\tau_{\varphi}^{R}$ and $\tau_{\varphi}^{E}$ are extracted from the measurements of Ramsey oscillations and Hahn echo measurements, respectively, as $\tau_{\varphi}^{R,E}=1/\gamma_{\varphi}^{R,E}=1/\left(\gamma_{2}^{R,E}-\gamma_{1}/2\right)$, where the decoherence rates $\gamma_{2}^{R,E}$ describe Ramsey oscillations and Hahn echo experiments [42, 39]. While in the low-temperature limit $\tau_{\varphi}^{R,E}\gtrapprox\tau_{1}$, with $\tau_{\varphi}^{R}$ exhibiting a clearly visible saturation below $50-70$ mK, the dephasing time starts decreasing at lower temperatures than $\tau_{1}$ and reaches $1-2\,\mu$s at 150 mK, becoming difficult to measure at $220-250\,$mK due to the finite ring-up time of the readout resonator $\sim Q_{\mathrm{ext}}/\omega_{r}\simeq 1\,\mu$s. Similarly to the analysis of the relaxation time, we can estimate the influence of quasiparticles on the dephasing, using the approach presented in Ref. [46]. The pure dephasing rate due to tunneling of equilibrium quasiparticles is given by expression
@@
The working temperature range is limited by the thermal quasiparticles in the basic operation as a cryostat thermometer. Yet when measuring objects on a chip coupled, e.g., by photon exchange, such a limitation becomes irrelevant, and the thermometer temperature range can be much wider (see further discussion in Section V.4).
V.2 Signal-to-Noise Ratio and Accuracy
Ultimately, the accuracy of the thermometer is limited by the quantum nature of the probe, which is reflected by the Cramer-Rao bound [70]. This limit can be obtained from the quantum Fisher information (QFI) derived for a particular probe [27]. For example, a two-level thermometer with energy separation $\hbar\omega_{ge}$ has the relative error of a single measurement $\left|\Delta T/\langle T\rangle\right|_{\mathrm{sm}}$ defined as
$\left(\frac{\Delta T}{\langle T\rangle}\right)^{2}_{\mathrm{sm}}\geq\frac{\left(1+e^{x_{ge}}\right)^{2}}{x_{ge}^{2}e^{x_{ge}}},$ (30)
$x_{ge}=\hbar\omega_{ge}/k_{B}T$. Analogously an expression for a three-level anharmonic probe can be obtained as
$\left(\frac{\Delta T}{\langle T\rangle}\right)^{2}_{\mathrm{sm}}\geq\frac{\left(e^{x_{ge}+x_{gf}}+e^{x_{ge}}+e^{x_{gf}}\right)^{2}}{\left[x_{ge}^{2}e^{x_{gf}}+x_{gf}^{2}e^{x_{ge}}+\left(x_{ge}-x_{gf}\right)^{2}\right]e^{x_{ge}+x_{gf}}},$ (31)
where $x_{gf}=\hbar\omega_{gf}/k_{B}T$ and $\omega_{gf}=\omega_{ge}+\omega_{ef}$. Taking into account averaging over $N_{\mathrm{av}}=2^{17}=131072$ repetitions, one gets the lower bound for the relative error $\left(\Delta T/\langle T\rangle\right)^{2}=\left(\Delta T/\langle T\rangle\right)^{2}_{\mathrm{sm}}/N_{\mathrm{av}}\geq 5.0\cdot 10^{-5}$ or $\left|\Delta T/\langle T\rangle\right|\geq 0.7\%$ for $\langle T\rangle=65$ mK and qubit frequency $\omega_{ge}=7.04$ GHz. The noise equivalent temperature $\mathrm{NET}=\sqrt{\left|\Delta T\right|^{2}\Delta t_{\mathrm{meas}}}$ with the measurement time $\Delta t_{\mathrm{meas}}=29$ s for our setup in the QFI-limited case would be $\mathrm{NET}_{\mathrm{QFI}}\approx 2.5$ mK$/\sqrt{\mathrm{Hz}}$. However, in practice such values can hardly be achieved due to poor SNR of the measurement system, unless amplifiers with quantum limited noise are used [42]. In our case, the quantum limit for the temperature error is one order of magnitude lower than the experimental errors. For example, Fig. 6 a represents a temperature time trace which was measured during several hours at the base cryostat temperature $T_{\mathrm{MXC}}=22$ mK. We calculate the average temperature and the deviation based on this statistics and recalculate it to the relative error per single measurement $\left|\Delta T/\langle T\rangle\right|_{\mathrm{sm}}=\sigma/\left(\langle T\rangle\sqrt{N_{\mathrm{av}}}\right)$. The data have the relative errors $\left|\Delta T_{A}/\langle T_{A}\rangle\right|=14-42\%$, $\left|\Delta T_{B}/\langle T_{B}\rangle\right|=8.6-9.4\%$ and $\left|\Delta T_{C}/\langle T_{C}\rangle\right|=8.7-9.4\%$, which corresponds to $\mathrm{NET}\approx 28.5$ mK$/\sqrt{\mathrm{Hz}}$ for $T_{B}$ and $T_{C}$. It can be seen, that for each sample all data points have overlapping confidence intervals of $2\sigma$ (see Fig. 6). Typically, the temperatures extracted from the population ratio $A$ tend to have a worse accuracy and underestimate the qubit effective temperature, compared to $B$ and $C$, which give close values of the average temperatures and deviations. Thus, one can simply use the measurement with the minimal error to derive the temperature.
Figure 6: a) Time traces of the effective qubit temperature for sample Q2-III at the base cryostat temperature $T_{\mathrm{MXC}}\approx 22\,$mK (shown as the black dashed line). b) Relative error of a single temperature measurement as a function of normalized effective temperature. The symbols denote experimental time averaged data obtained from time traces of three samples: R2-I (green), R3-II (blue) and Q2-III (red). The error bars show the standard deviation. We also show data from other works: Rabi oscillation measurements [17] (grey stars), correlations [18] (open and solid black circles) and original paper, where the $\pi$-pulse technique was described [20] (grey triangle and diamond). The curves show the minimal achievable error for each experiment, which is defined by the Cramer-Rao bound and QFI, i.e. all data points should be above it. Solid red curve is the QFI limit obtained from Eq. (31) for three-level case and the black dashed line depicts the two-level case Eq. (30). More information on the data shown in the figure can be found in Tab. 4.
Accuracy of this thermometry technique can be compared with other works [20, 18]. In the latter a quantum limited JPA with SNR $\sim 6$ was used for the measurements, which had shown a high relative accuracy of determining the excited state population ($\Delta p_{e}/p_{e}\approx 5\%$). In order to compare different measurement techniques, we calculated single measurement errors with respect to $N_{\mathrm{av}}$ for each experiment (which we take as standard deviation $\Delta T=\sigma$) and NET values for each experiment, using the data presented in the papers. We also argue, that the lowest effective qubit temperature shown in Ref. [18] was calculated correctly, and recalculate it from the excited state population using Eq. (3). Noteworthy, if the effective qubit temperature obtained in Ref. [18] had indeed matched the MXC stage temperature of 22 mK, the accuracy per single measurement would have significantly exceeded the Cramer-Rao limit. Results are shown in Fig. 6 b and in Tab. 4. Accuracy shown by the correlation measurements using a JPA almost reach the QFI limit, whereas other measurements appeared to be comparable to our results, which demonstrates the effectiveness of quantum limited amplifiers in improving the SNR.
Table 4: Comparison with other experiments on qubit thermometry. For sample Q2-III three best measurements for $A$, $B$ and $C$ measurements are shown. Easiest comparison can be done by looking at error per single measurement $\left|\frac{\Delta T_{\mathrm{eff}}}{T_{\mathrm{eff}}}\right|_{\mathrm{sm}}$ or at the noise equivalent temperature (NET).
@@ V.2 Signal-to-Noise Ratio and Accuracy
Sample ID Measurement $T_{\mathrm{eff}}$, mK $\omega_{ge}/2\pi$, GHz $N_{\mathrm{av}}$ $\left|\frac{\Delta T_{\mathrm{eff}}}{T_{\mathrm{eff}}}\right|_{\mathrm{sm}}$ $\Delta t_{\mathrm{meas}}$, s NET, mK$/\sqrt{\mathrm{Hz}}$ $T_{\mathrm{MXC}}$, mK
Q2-III $A2$ $60.1\pm 8.3$ 7.04 $2^{17}$ 49.8 29 44.7 22
$B1$ $61.3\pm 5.3$ 7.04 $2^{17}$ 31.4 2.84 .5 22
$C1$ $61.3\pm 5.3$ 7.04 $2^{17}$ 31.3 2.84 .5 22
Ref. [17] lowest $T_{\mathrm{MXC}}$ $31.8\pm 1.42$ 4.97 $1.5\cdot 10^{6}$ 20.8 10 43.2 15
highest $T_{\mathrm{MXC}}$ $59.8\pm 0.44$ 4.97 $1.5\cdot 10^{6}$ 15.6 10 31.2 60
Ref. [18] JPA off $52.63\pm 3.11$ 6.26 $2^{20}$ 91.0 900 93.3 22
JPA on $52.63\pm 0.15$ 6.26 $2^{20}$ 4.30 900 4.45 22
Ref. [20] $A$ $169.9\pm 14.4$ 6.74 $6\cdot 10^{4}$ 20.8 9 43.2 13
$B,C$ $162.8\pm 10.4$ 6.74 $6\cdot 10^{4}$ 15.6 9 31.2 13
Notably, if a thermometer has a constant level separation, the relative error grows exponentially at $T<\hbar\omega_{ge}/2k_{B}$ (see Fig. 6), which is, demonstrated to some extent by experimental data from Ref. [17]. Our system is a three-level quantum thermometer with a constant anharmonic spectrum (“sub-optimal quantum thermometers” [27]). Knowing the dependence of QFI on the spectrum of the system, it is possible to optimize it for a better performance in the experimentally realistic temperature range by choosing an appropriate level separation and anharmonicity. A qubit with a flux-tunable spectrum could provide higher precision for thermometry, by appropriate adjustment of the ratio $\hbar\omega_{ge}/k_{B}T_{\mathrm{eff}}$.
@@
Other parameters of the experimental setup can also be fine-tuned to improve both the temperature measurement accuracy and the range of the cryostat temperatures, so that it reaches the quasiparticle limit. First of all, improvement of the $\pi$-pulse efficiency affects positively on the measurement accuracy. Secondly, as mentioned above, the short lifetime of the qubit is the main restriction of the method. We can divide the qubit state decay into two parts: i) decay during the control sequence and ii) decay during the readout. The first process leads to change of the qubit population distribution $p_{i}$ and can be associated to imperfect control pulses, leading to increased deviation between $T_{A1},T_{A2},\dots,T_{C3}$. Decay during the readout pulse is defined mainly by the ring-up time of the resonator, and diminishes distinguishability between the qubit states (i.e. difference between $\varphi_{i}$ on the $I-Q$ plane), thus reducing the SNR, and also distorting the extracted effective temperature. The main method of improving the SNR, except for using a quantum limited amplifier, is optimizing the ratio $\chi/\kappa$ for the readout resonator, where $\chi$ is the dispersive shift and $\kappa$ is the linewidth, as well as an appropriate readout frequency detuning for a given $\chi/\kappa$, which directly affects the pure state responses $\varphi_{i}$ [71].
@@
V.3 Operation Speed